OpenStudy (anonymous):

HELP PLEASEE!! I NEED ITI DONT INDERSTAND HOW TO FIGURE THIS OUT HELPP!!! Jeff and Lucy have been asked to wash their mom’s minivan. It takes Jeff 2 hours to wash the van by himself, and it takes Lucy 1.5 hours to wash the van by herself. How long will it take Jeff and Lucy to wash the van if they work together? a.What equation is used to solve this problem? What does each variable represent? b.Identify the given information. c.Enter the given information into the equation. d.Solve the equation. e.State the solution.

OpenStudy (anonymous):

wow a through e, but you can do it in one step

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

it will take them \[\frac{2\times 1.5}{2+1.5}\] hours

OpenStudy (anonymous):

i mean you can write down what you need to compute in one step you still have to compute it

OpenStudy (anonymous):

\[\frac{2\times 1.5}{2+1.5}=\frac{3}{3.5}=\frac{30}{35}=\frac{6}{7}\]

OpenStudy (anonymous):

we can work through a to e if you like, make your teacher happy

OpenStudy (anonymous):

I STILL DONT UNDERSTAND HELPPLZZ

OpenStudy (anonymous):

jeff takes two hour to wash the van, which means he washes one half of the van per hour that his, his rate is \(\frac{1}{2}\)

OpenStudy (anonymous):

lucy takes \(1.5\) hours, which means her rate per hour is \(\frac{1}{1.5}=\frac{2}{3}\) in other words she washes two thirds of the van per hour

OpenStudy (anonymous):

to find their combined rate, the rate when they work together, add the two rates you get their combined rate is \[\frac{1}{2}+\frac{2}{3}=\frac{3+4}{3\times 2}=\frac{7}{6}\]

OpenStudy (anonymous):

you want to know how long it takes for them to wash one van, so set \[\frac{7}{6}t=1\] and solve for \(t\) via \[t=\frac{6}{7}\]